A plane can fly at a 1720 miles in 5 hours against the wind at max speed. The same plane can fly 1720 miles in 4 hours with the wind at max speed. What is the speed of the plane with no wind? Let x = max speed, and y = wind speed.
Answer by rfer(12647)

x-y=430
x+y=344
(x=344-y)
344-y-y=430
-2y=86
-y=43
y=-43
wind speed =43mph
airplane speed =387mph